OpenStudy (anonymous):

If cos θ=–5/13 and sin θ < 0, what is tan θ ? A. 12/5 B. 13/5 C. -12/5 D. sq. rt 194/5

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I think it's C...

OpenStudy (anonymous):

sorry, use cos = adj/hyp

OpenStudy (anonymous):

tan = opp/adj

OpenStudy (anonymous):

so make the triangle and solve for the missing side

OpenStudy (anonymous):

is it c?

OpenStudy (anonymous):



OpenStudy (anonymous):

find the missing side using Pythagorean theorem

OpenStudy (anonymous):

those are squared, correct?

OpenStudy (anonymous):

13^2 = 5^2 + x^2

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

yep, now find tan which is opp/adj

OpenStudy (anonymous):